Two balls are drawn at random from an urn containing 5 white and 3 red balls. The first ball is not replaced before the second ball is drawn. Find the probability that the first ball is red, conditional on the second ball being red. 
 
i am getting 2/7. can someone pls clarify 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
required probability = (3/8 * 2/7) /[(3/8 * 2/7) + (5/8* 3/7)]
 
 = 6/21 = 2/7 is correct answer
 
Regards
Arun